# Randomized methods for dynamical low-rank approximation

### v2

@@ 4.4 Kinetic equations with Vlasov-Poisson
[…]
where $f(t,x,v)$ is the particle-density function with $x\in\Omega_{x}\subset\mathbb{R}^{d}$ and $v\in\Omega_{v}\subset\mathbb{R}^{d}$. One aspect of equation (27) is its large memory footprint since it requires to store $O(n^{2d})$ floating point numbers, where $n$ is the number of grid points. To overcome the memory limitations, low-rank techniques have been introduced in [EL18] and we refer to references therein for state-of-the-art full rank solvers. The research on low-rank techniques for the Vlasov equation is still in development, see e.g. [GQ22, CE22, GQ24, UZ24]. Note that in the literature, the discretization is typically done after deriving equations for the low-rank factors of the function $f(t,x,v)$ itself. In this paper, we first perform a spatial discretization via finite differences and get the following matrix differential equation:
@@
[…]
@@
The Landau damping is a phenomenon that occurs in plasma and can be modeled with equation (27). It is a damping of the electric field due to the interaction of the particles with the field. We consider the same setup as provided in [EL18], that is, the domain is $\Omega=(0,4\pi)\times(-6,6)$ with periodic boundary conditions and initial value
@@
[…]

### v3

@@ 4.4 Kinetic equations with Vlasov-Poisson
$\nabla\cdot E(f)(t,x)=-\int f(t,x,v)dv+1,\qquad\nabla\times E(f)(t,x)=0,$
where $f(t,x,v)$ is the particle-density function with $x\in\mathcal{D}_{x}\subset\mathbb{R}^{d}$ and $v\in\mathcal{D}_{v}\subset\mathbb{R}^{d}$. One aspect of equation (27) is its large memory footprint since it requires to store $O(n^{2d})$ floating point numbers, where $n$ is the number of grid points. To overcome the memory limitations, low-rank techniques have been introduced in [EL18] and we refer to references therein for state-of-the-art full rank solvers. The research on low-rank techniques for the Vlasov equation is still in development, see e.g. [GQ22, CE22, GQ24, UZ24]. Note that in the literature, the discretization is typically done after deriving equations for the low-rank factors of the function $f(t,x,v)$ itself. In this paper, we first perform a spatial discretization via finite differences and get the following matrix differential equation:
@@
where $D_{x}$ and $D_{v}$ are (high-order) centered finite difference matrices, $V$ and $E(t,A(t))$ are two diagonal matrices with coefficients derived from (27). In the following experiments, the reference solver for solving equation (28) is scipy RK45 solver with tolerance $10^{-8}$.
@@
The Landau damping is a phenomenon that occurs in plasma and which can be modeled with equation (27). It is a damping of the electric field due to the interaction of the particles with the field. We consider the same setup as provided in [EL18], that is, the domain is $\Omega=(0,4\pi)\times(-6,6)$ with periodic boundary conditions and initial value
@@
Figure 5: Numerical simulations of the linear Landau damping (29) with stepsize $h=0.04$. The projector-splitting is performed with a Strang splitting ($\mathrm{KSL2}$) and the substeps are solved with the reference solver. The DRSVD is performed with oversampling parameter $p=5$ and $q=1$ power iteration. The DGN is performed with oversampling parameters $p=5$, $\ell=0$ and $q=1$ power iteration.